OpenStudy (anonymous):

2x – y = 5 3x – 2y = 12

jimthompson5910 (jim_thompson5910):

2x – y = 5 – y = -2x + 5 y = 2x - 5 -------------------- 3x – 2y = 12 3x – 2(2x - 5) = 12 3x-4x+10=12 -x+10=12 -x=12-10 -x=2 x=2/(-1) x=-2 -------------------------- y = 2x - 5 y = 2(-2) - 5 y = -4 - 5 y = -9 So the solutions are x = -2 and y = -9